# Disentangled Representation Learning through Geometry Preservation with the Gromov-Monge Gap

### v1

@@ 3.3 Estimation and Computation from Samples
Computing $\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$ requires solving a discrete GW problem between $r_{n}$ and $T\sharp r_{n}$ to obtain $\mathrm{GW}_{c_{\mathcal{X}},c_{\mathcal{Y}}}({r}_{n},T\sharp{r}_{n})$. To alleviate computational challenges, we estimate this term using an entropic regularization $\varepsilon\geq 0$, as introduced in Eq. (EGWP):
$\mathcal{GM}_{{r}_{n},\varepsilon}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)\vcentcolon=\mathcal{D}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)-\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n},\varepsilon}({r}_{n},T\sharp{r}_{n})\,.$ (4)
Choosing $\varepsilon=0$, we recover the unregularized one $\mathcal{GM}_{{r}_{n},0}^{c_{\mathcal{X}},c_{\mathcal{Y}}}=\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}$. Moreover, the entropic estimator preserves the positivity, as for $\varepsilon\geq 0$, we have $\mathcal{GM}_{{r}_{n},\varepsilon}^{c_{\mathcal{X}},c_{\mathcal{Y}}}\geq 0$ (see A.1). As described in § 2, we compute $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{{r}_{n},\varepsilon}({r}_{n},T\sharp{r}_{n})$ using Peyré et al. [54]’s solver. While it always has $\mathcal{O}(n^{2})$ memory complexity, when $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$ or $c_{\mathcal{X}}=c_{\mathcal{Y}}=\|\cdot-\cdot\|_{2}^{2}$, this solver runs in $\mathcal{O}(n^{2}d)$ time [62, Alg. 2]. Since the cosine similarity is equivalent to the inner product, up to pre-normalization of $\mathbf{x}_{i}$ and $T(\mathbf{x}_{i})$, the computation of the GMG for the costs of interest [i] or [ii] scales as $\mathcal{O}(n^{2}d)$ in time. We use ott-jax’s [16] implementation of this scheme.
@@
As laid out, the GMG can be used as a regularization loss to push any model $T$ to be more geometry-preserving. A natural question that arises when defining such a regularizer is: what are its regularity properties, and in particular, is it convex? In the following, we study the convexity of $T\mapsto\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$, and its finite-sample counterpart $T\mapsto\mathcal{GM}_{{r}_{n}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T)$. We focus on the costs of interest [i] or [ii]. For simplicity, we replace cosine similarity with inner product—i.e., $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$—as they are equivalent, up to normalization of $r$ and $T$. We then study the convexity of the GMG for (i) the (scaled) squared Euclidean distances and (ii) the inner product, denoted respectively by (i) $\mathcal{GM}_{r}^{2}$ and (ii) $\mathcal{GM}_{r}^{\langle\cdot,\cdot\rangle}$. To that end, we introduce a weaker notion of convexity, previously defined for functions on $\mathbb{R}^{d}$ [17], which we extend here to $L_{2}(r)=\{T\,|\,\|T\|_{L_{2}(r)}^{2}\vcentcolon=\int_{\mathcal{X}}\|T(\mathbf{x})\|_{2}^{2}\mathop{}\!\mathrm{d}r(\mathbf{x})<+\infty\}$.
@@
Both $\mathcal{GM}_{r}^{2}$ and $\mathcal{GM}_{r}^{\langle\cdot,\cdot\rangle}$, as well as their finite sample versions, are weakly convex.
  • Finite sample. We note $\mathbf{X}\in\mathbb{R}^{n\times d}$ the matrix that stores the $\mathbf{x}_{i}$, i.e. the support of $r_{n}$, as rows. Then, (i) $\mathcal{GM}_{r_{n}}^{2}$ and (ii) $\mathcal{GM}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ are respectively (i) $\gamma_{2,n}$ and (ii) $\gamma_{\textrm{inner},n}$-weakly convex, where: $\gamma_{\textrm{inner},n}=\lambda_{\max}(\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top})$ and $\gamma_{2,n}=\gamma_{\textrm{inner},n}+\max_{i=1\dots n}\|\mathbf{x}_{i}\|_{2}^{2}$.
  • Asymptotic. (i) $\mathcal{GM}_{r}^{2}$ and (ii) $\mathcal{GM}_{r}^{\langle\cdot,\cdot\rangle}$ are respectively (i) $\gamma_{2}$ and (ii) $\gamma_{\textrm{inner}}$-weakly convex, where: $\gamma_{\textrm{inner}}=\lambda_{\max}(\mathbb{E}_{\mathbf{x}\sim r}[\mathbf{x}\mathbf{x}^{\top}])$ and $\gamma_{2,n}=\gamma_{\textrm{inner}}+\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$.
@@
  • First, we have $\gamma_{2}\geq\gamma_{\textrm{inner}}$. Therefore, $\mathcal{GM}_{r}^{2}$ is less convex than $\mathcal{GM}_{r}^{\langle\cdot,\cdot\rangle}$, making it harder to optimize, and the same argument holds for their estimator. In other words, we provably recover that, in practice, preserving the (scaled) distances is harder than simply preserving the angles.
@@
  • Third, and probably the most surprising, the finite sample GMG is more convex in high dimension. Indeed, $\gamma_{\textrm{inner},n}$ is the spectral width of $\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top}$, containing the (rescaled) inner-products between the $\mathbf{x}_{i}\sim r$. When $n>d$, $\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top})=0$ as $\mathrm{rank}(\mathbf{X}\mathbf{X}^{\top})=d$. Then, $\gamma_{\textrm{inner},n}$ increases, which in turn decreases the GMG’s convexity. On the other hand, when $d>n$, $\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top})>0$ if $\mathbf{X}$ is full rank. Intuitively, $\mathcal{GM}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ is nearly convex when $\mathbf{X}\mathbf{X}^{\top}$ is well conditioned. Assuming that the $\mathbf{x}_{i}$ are normalized, this might happen in high dimension, as those points will be orthogonal with high probability. This property suggests that, in practice, and contrary to the insights provided by the statistical OT literature [71, 25, 56, 73], the GMG might not benefit a large sample size.
@@ 3.5 Learning with the Gromov-Monge gap
Provided a source distribution $p$ and a target $q$ defining a marginal constraint, learning with the GMG remains to optimize a loss of the form
$\mathcal{L}(\theta)\vcentcolon=\Delta(T_{\theta},p,q)+\lambda_{\textrm{GMG}}\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(T_{\theta})$ (5)
where $\Delta$ is a fitting loss, which can access paired, or unpaired, samples of $p$ and $q$. In theory, from Prop. (3.3), we can choose any reference $r$ s.t. $\mathrm{Spt}(p)\subset\mathrm{Spt}(r)$. In practice, given the insights of Thm. (3.6), we usually consider $r=p$. Note that replacing $\mathcal{GM}_{r}^{c_{\mathcal{X}},c_{\mathcal{Y}}}$ by $\mathcal{D}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{r}$ in Eq. (5), we similarly define the learning procedure with the distortion. We compare their effect in Figure 1.
@@
In the VAE setting, (i) when we apply the GMG (or the distortion) to the encoder $e_{\phi}$, the fitting loss is defined through the prior matching constraint, as described in § 2.1. Conversely, (ii) when we apply the GMG to the decoder $d_{\phi}$, the fitting loss is defined through the reconstruction loss. Additionally, in both cases, our goal is to promote the latent space to preserve certain geometric features of the data. Therefore, in (i) we use $r=p_{\textrm{data}}$ the data distribution as reference $r$, while in (ii) we use the latent distribution $r=q_{\phi}$. Introducing weightings $\lambda_{\textrm{enc}},\lambda_{\textrm{dec}}\geq 0$, determining which mapping we regularize, this remains to optimize the loss
$\mathcal{L}_{\beta\textrm{-GMG}}(\theta,\phi)=\mathcal{L}_{\beta}(\theta,\phi)+\lambda_{\textrm{enc}}\mathcal{GM}_{p_{\textrm{data}}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(e_{\phi})+\lambda_{\textrm{dec}}\mathcal{GM}_{q_{\phi}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(d_{\theta}),$ (6)
where $\mathcal{L}_{\beta}$ is introduce in § 2.1. Note that this loss can easily be extended to $\beta$-TCVAE and the combination of other regularization terms. While previous work [51, 41] chooses to apply the geometric regularizations to the decoder, we investigate regularizing both, separately or simultaneously. For completeness, we also derive the VAE-loss when learning with the distortion:
$\mathcal{L}_{\beta\textrm{-DST}}(\theta,\phi)=\mathcal{L}_{\beta}(\theta,\phi)+\lambda_{\textrm{enc}}\mathcal{D}_{p_{\textrm{data}}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(e_{\phi})+\lambda_{\textrm{dec}}\mathcal{D}_{q_{\phi}}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(d_{\theta}),$ (7)
The choice of $c_{\mathcal{X}},c_{\mathcal{Y}}$.
@@ 4 Experiments
Table 1: Effect of different geometric regularization on disentanglement (DCI-D, Shapes3D [33]). We highlight the best method per regularization, and the best/second best per column.
@@
Isometric (IR)
+ (DST) 71.5 ±$\pm 3.6$ 75.8 ±$\pm 6.6$ 92.1 ±$\pm 9.7$ 90.9 ±$\pm 7.6$
+ (GMG) 72.0 ±$\pm 12.5$ 78.9 ±$\pm 5.0$ 92.5 ±$\pm 4.4$ 91.7 ±$\pm 6.0$
Scaled Isometric (SIR)
+ Jacobian 61.4 ±$\pm 12.8$ 76.7 ±$\pm 4.5$ 90.5 ±$\pm 3.8$ 91.5 ±$\pm 5.6$
+ (DST) 67.4 ±$\pm 7.1$ 77.9 ±$\pm 4.5$ 93.2 ±$\pm 9.7$ 94.5 ±$\pm 6.9$
+ (GMG) 70.0 ±$\pm 5.9$ 81.0 ±$\pm 3.2$ 93.3 ±$\pm 8.6$ 96.1 ±$\pm 3.8$
Conformal (CR)
+ (DST) 76.8 ±$\pm 4.1$ 81.3 ±$\pm 4.7$ $87.5$$\pm 3.3$ $91.9$$\pm 9.4$
+ (GMG) 82.1 ±$\pm 4.5$ 83.7 ±$\pm 8.8$ 95.7 ±$\pm 5.8$ 96.9 ±$\pm 4.9$
Experimental setup.
We evaluate the effectiveness of the (GMG) as regularizer in disentangled representation learning. We use the standard $\beta$-VAE and $\beta$-TCVAE as our base models and incorporate the GMG on top of them. Moreover, we consider the recently proposed HFS [60] regularization on top of both $\beta$-VAE and $\beta$-TCVAE, totaling four base models. Our primary goal is to investigate the differences between using the GMG and the (DST) as regularizers, specifically examining whether the (GMG) leads to more disentangled representations compared to the raw distortion. Additionally, we aim to determine which geometric regularization (IR, SIR, CR) is most beneficial for disentanglement and what part of the pipeline should be regularized. Lastly, we investigate whether a geometric regularization can help prevent the collapse of learned representation in the Decoder-free setting. We evaluate the learned latents with DCI-D [20] as it was found to be the metric most suitable to measure disentanglement [44, 18]. We benchmark over multiple datasets commonly used to assess the performances of disentangled representation learning methods: Shapes3D [33], DSprites [27], SmallNORB [40], and Cars3D [57].
4.1 Evaluating Different Geometric Regularizations
Regularizing the Decoder.
First, we focus on the difference between optimizing for different geometry-preserving regularizations. We compare between IR, SIR, and CR [41] realized through either the (DST), or (GMG). Additionally, we include the Jacobian-based SIR as introduced in Lee et al. [41]. We report full results on Shapes3D [9] over 5 seeds in Table 1. We observe that the GMG outperforms the sole distortion loss for all levels of regularization and baselines. Moreover, we find that a CR performs best with respect to disentanglement compared to both IR and SIR. Note, that employing a CR has not been benchmarked for disentangled representation learning before. These results elucidate the clear benefit of using the GMG in its CR implementation in terms of learning more disentangled representations significantly improving upon previously proposed regularization.
@@ 4.1 Evaluating Different Geometric Regularizations
Table 2: Effect of (GMG) and (DST) leveraged as a conformal regularization (CR) on the disentanglement of learned representations as measured by DCI-D over four datasets. We highlight the best, and second best result for each dataset and method.
CR $\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
Shapes3D [33]
Base $65.8$$\pm 15.6$ $75.0$$\pm 3.4$ $88.1$$\pm 7.4$ $90.2$$\pm 7.5$
+ (DST) 76.8 ±$\pm 4.1$ 81.3 ±$\pm 4.7$ $87.5$$\pm 3.3$ $91.9$$\pm 9.4$
+ (GMG) 82.1 ±$\pm 4.5$ 83.7 ±$\pm 8.8$ 95.7 ±$\pm 5.8$ 96.9 ±$\pm 4.9$
DSprites [27]

### v2

@@ 3.3 Estimation and Computation from Samples
Computing the GMG requires solving a discrete GW problem between $r_{n}$ and $T\sharp r_{n}$ to get $\mathrm{GW}({r}_{n},T\sharp{r}_{n})$. We compute this term using an entropic regularization $\varepsilon\geq 0$, as in Eq. (EGWP):
$\mathrm{GMG}_{{r}_{n},\varepsilon}(T)\vcentcolon=\mathrm{DST}_{{r}_{n}}(T)$ (4)
$\,\,-\mathrm{GW}_{\varepsilon}({r}_{n},T\sharp{r}_{n})\,.$
Choosing $\varepsilon=0$, we recover $\mathrm{GMG}_{{r}_{n},0}(T)=\mathrm{GMG}_{{r}_{n}}(T)$. Moreover, the entropic estimator preserves positivity, as for $\varepsilon\geq 0$, $\mathrm{GMG}_{r_{n},\varepsilon}(T)\geq 0$ (see B.3). We compute $\mathrm{GW}_{\varepsilon}({r}_{n},T\sharp{r}_{n})$ using Peyré et al. (2016)’s solver introduced in 2.2. We use the implementation provided by ott-jax (Cuturi et al., 2022). In practice, we select $\varepsilon$ based on (positive) statistics from the cost matrices $\mathbf{C}_{\mathcal{X}}$,$\mathbf{C}_{\mathcal{Y}}$. We define a scale $\varepsilon_{0}$ and set $\varepsilon=\varepsilon_{0}\cdot\texttt{stat}(\mathbf{C}_{\mathcal{X}})\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$. Standard options for the statistic include $\texttt{stat}\in\{\texttt{mean},\texttt{max},\texttt{std}\}$. This procedure is equivalent to running the entropic GW solver on the re-scaled cost matrices $\mathbf{C}_{\mathcal{X}}/\texttt{stat}(\mathbf{C}_{\mathcal{X}})$ and $\mathbf{C}_{\mathcal{Y}}/\texttt{stat}(\mathbf{C}_{\mathcal{Y}})$ with $\varepsilon=\varepsilon_{0}$; see App. B.2. It is a common practical trick, initially suggested for stabilizing the Sinkhorn algorithm Cuturi (2013), then extended to the entropic GW solver (Cuturi et al., 2022). Additionally, as suggested by Scetbon et al. (2022), to promote the solver’s stability, we always initialize it with $\mathbf{P}_{\textrm{init}}=\tfrac{1}{n^{2}}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$.
Computational Complexity.
For usual costs, including inner products, $\ell_{p}^{q}$ distances, and standard CPD kernels, the computation of $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})$ (resp. $c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})$) can be done in $\mathcal{O}(d_{\mathcal{X}})$ time (resp. $\mathcal{O}(d_{\mathcal{Y}})$). Consequently, the DST can be computed in $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$ time. Furthermore, as discussed in § 2.2, the time complexity of the entropic GW solver is $\mathcal{O}(n^{3})$ in general, but can be reduced to $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$ when $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$, or $c_{\mathcal{X}}=\|\cdot-\cdot\|_{2}^{2},c_{\mathcal{Y}}=\alpha\|\cdot-\cdot\|_{2}^{2}$. Therefore, since the cosine similarity is equivalent to the inner product, up to pre-normalization of $\mathbf{x}_{i}$ and $T(\mathbf{x}_{i})$, this solver runs in $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$ for the costs of interest $\mathbf{(Sc)L2^{2}}$ and $\mathbf{Cos}$. The complete algorithm, along with a time complexity analysis of each step, is described in Alg. 1.
@@
[…]
@@
Both $\mathrm{GMG}_{r}^{2}$ and $\mathrm{GMG}_{r}^{\langle\cdot,\cdot\rangle}$, as well as their finite sample versions, are weakly convex.
  • Finite sample. We note $\mathbf{X}\in\mathbb{R}^{n\times d}$ the matrix that stores the $\mathbf{x}_{i}$, i.e. the support of $r_{n}$, as rows. Then, (i) $\mathrm{GMG}_{r_{n}}^{2}$ and (ii) $\mathrm{GMG}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ are respectively (i) $\gamma_{2,n}$ and (ii) $\gamma_{\textrm{inner},n}$-weakly convex, where: $\gamma_{\textrm{inner},n}=\lambda_{\max}(\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top})-\lambda_{\min}(\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top})$ and $\gamma_{2,n}=\gamma_{\textrm{inner},n}+\max_{i=1\dots n}\|\mathbf{x}_{i}\|_{2}^{2}$.
  • Asymptotic. (i) $\mathrm{GMG}_{r}^{2}$ and (ii) $\mathrm{GMG}_{r}^{\langle\cdot,\cdot\rangle}$ are respectively (i) $\gamma_{2}$ and (ii) $\gamma_{\textrm{inner}}$-weakly convex, where: $\gamma_{\textrm{inner}}=\lambda_{\max}(\mathbb{E}_{\mathbf{x}\sim r}[\mathbf{x}\mathbf{x}^{\top}])$ and $\gamma_{2,n}=\gamma_{\textrm{inner}}+\max_{\mathbf{x}\in\mathrm{Spt}(r)}\|\mathbf{x}\|_{2}^{2}$.
@@
  • First, we have $\gamma_{2}\geq\gamma_{\textrm{inner}}$. Therefore, $\mathrm{GMG}_{r}^{2}$ is less convex than $\mathrm{GMG}_{r}^{\langle\cdot,\cdot\rangle}$, making it harder to optimize, and the same argument holds for their estimator. In other words, we provably recover that, in practice, preserving the (scaled) distances is harder than simply preserving the angles.
@@
  • Third, and most surprising, the finite sample GMG is more convex in high dimension. Indeed, $\gamma_{\textrm{inner},n}$ is the spectral width of $\tfrac{1}{n}\mathbf{X}\mathbf{X}^{\top}$, which contains the (rescaled) inner products between $\mathbf{x}_{i}\sim r$. When $n>d$, $\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top})=0$ as $\mathrm{rank}(\mathbf{X}\mathbf{X}^{\top})=d$. Then, $\gamma_{\textrm{inner},n}$ increases, which in turn decreases the GMG’s convexity. However, when $d>n$, $\lambda_{\min}(\mathbf{X}\mathbf{X}^{\top})>0$ if $\mathbf{X}$ is full rank. Intuitively, $\mathrm{GMG}_{r_{n}}^{\langle\cdot,\cdot\rangle}$ is nearly convex when $\mathbf{X}\mathbf{X}^{\top}$ is well conditioned. Assuming that the $\mathbf{x}_{i}$ are normalized, this might happen in high dimension, as they will be orthogonal with high probability. This suggests that, contrary to the insights provided by the statistical OT literature (Weed and Bach, 2017, Zhang et al., 2023), the GMG might not benefit a large sample size.
@@ 3.5 Learning with the Gromov-Monge gap
Given a source $p$ and a target distribution $q$, we can use the GMG to guide a parameterized map $T_{\theta}$ towards approximating a Gromov-Monge map between $p$ and $q$. We handle the marginal constraint $T_{\theta}\sharp p=q$ separately through a fitting loss $\Delta(T_{\theta},p,q)$. Provided any reference $r$ s.t. $\mathrm{Spt}(p)\subset\mathrm{Spt}(r)$, we minimize
$\min_{\theta}\Delta(T_{\theta},p,q)+\lambda\mathrm{GMG}_{r}(T_{\theta})$ (5)
$\Delta$ can operate on paired (e.g., in VAE, the reconstruction loss), or unpaired (e.g., in VAE, the KL loss in VAE), samples of $p$ and $q$. Note that, in theory and as stated in Prop. (3.3), we can select any reference $r$ such that $\mathrm{Spt}(p)\subset\mathrm{Spt}(r)$. However, based on the insights from Thm. (3.4), we typically choose $r$ with minimal support size and, in practice, set $r=p$. This learning procedure is illustrated in Fig. 2, where we also explore the effect of replacing $\mathrm{GMG}_{r}(T_{\theta})$ by $\mathrm{DST}_{r}(T_{\theta})$ in Eq. (5).
@@
In the VAE setting, we can use the GMG promote the (i) encoder $e_{\phi}$ or the (ii) decoder $d_{\theta}$ to mimic a Gromov-Monge map. In (i) we use $r_{e}=p_{\textrm{data}}$ the data distribution as reference $r$, while in (ii) we use the latent distribution $r_{d}=e_{\phi}\sharp p_{\textrm{data}}$. Introducing weightings $\lambda_{e},\lambda_{d}\geq 0$, determining which mapping we regularize, this remains to minimize
$\min_{\theta,\phi}\mathbb{E}_{\mathbf{x}\sim p_{\textrm{data}},\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}[\underbrace{-\log p_{\theta}(\mathbf{x}|\mathbf{z})}_{\textrm{(i) reconstruction}}+\underbrace{\beta D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))}_{\textrm{(ii) prior matching}}]+\underbrace{\lambda_{e}\mathrm{GMG}_{r_{e}}(e_{\phi})+\lambda_{d}\mathrm{GMG}_{r_{d}}(d_{\theta})}_{\textrm{(iii) geom. preservation}},$ (6)
With this loss, prior matching and geometric desiderata can be efficiently combined, as terms (ii) and (iii) can simultaneously be 0. Note that this loss can be easily extended to more advanced prior matching objectives, such as the $\beta$-TCVAE loss (Chen et al., 2018), and can be combined with other regularizers, such as the HFS (Roth et al., 2023). We explore this strategy in experiments § 4.2.
@@ 4 Experiments
4.1 GMG vs. DST: Illustrative Example
In Fig. 2, we start by illustrating the difference between the DST and GMG on synthetic data. We train $T_{\theta}$ to map from a $3$D source $p$ to a $2$D target distribution $q$ (first and second column) by minimizing $\mathcal{L}(\theta)\vcentcolon=\mathrm{S}_{\varepsilon}(T_{\theta}\sharp p,q)+\mathrm{R}(T_{\theta})$, where $\mathrm{S}_{\varepsilon}$ is the Sinkhorn divergence (Feydy et al., 2019) and $\mathrm{R}$ a geometry-preserving regularizer. We compare three settings, [i] no regularization $\mathrm{R}=0$, [ii] $\mathrm{R}=\mathrm{DST}_{p}$ as regularizer , and [iii] $\mathrm{R}=\mathrm{GMG}_{p}$. For all cases, we plot the transported distribution $T_{\theta}\sharp p$ after training. Without regularization (third column), we fit the marginal constraint $T_{\theta}\sharp p=q$ but do not preserve the geometric features. With the DST (fourth column), we preserve geometric features but do not fit the marginal constraint, i.e., $T_{\theta}\sharp p\neq q$. On the other hand, with the GMG (fifth column), we get the best compromise by approximating a Gromov-Monge map: we fit $T_{\theta}\sharp p=q$, while preserving the geometric features as fully as possible.
4.2 Leveraging the Gromov-Monge Gap for Disentanglement
Experimental Setup.  Having demonstrated how the DST enables (i) fitting a marginal constraint while (ii) preserving geometric features as much as possible, we now apply it to disentangled representation learning. Our primary goal is to investigate whether the GMG results in enhanced disentanglement compared to the DST by efficiently combining (i) prior matching with (ii) geometric constraints on the representation space. Moreover, we aim to determine which costs $c_{\mathcal{X}},c_{\mathcal{Y}}$ to choose in the geometric regularization and what part of the pipeline should be regularized, the encoder $e_{\phi}$, or the decoder $d_{\theta}$. The experimental setup is described below.
  • Baselines. We use the standard $\beta$-VAE and $\beta$-TCVAE as our starting models, with the option to apply the recent HFS regularization  (Roth et al., 2023) to each, resulting in a total of four base configurations. Note that the latter does not leverage geometric constraints; it is only used to enhance prior matching. We then investigate the effect of various geometry-preserving regularizations on disentanglement on top of these four base configurations. We consider GMG, DST and the Jacobian-based (Jac) regularization (Lee et al., 2022) discussed in 2.1. Given the inclusion of the DST, we naturally consider Nakagawa et al. (2023) as a baseline.
  • Metrics. We evaluate the learned representations using the DCI-D (Eastwood and Williams, 2018) as it was found that it is the metric most suitable for measuring the disentanglement (Locatello et al., 2020, Dittadi et al., 2021). We report mean and standard deviation over 5 seeds.
  • Datasets. We benchmark over four standard datasets: Shapes3D (Kim and Mnih, 2018), DSprites Higgins et al. (2017), SmallNORB LeCun et al. (2004), and Cars3D Reed et al. (2015).
[…]
Table 1: Effect of different regularization on disentanglement (DCI-D on Shapes3D). We highlight the best method per regularization type ($\mathbf{L2^{2}}$, $\mathbf{ScL2^{2}}$, or $\mathbf{Cos}$), and the best/second best per column.
@@ 4.2 Leveraging the Gromov-Monge Gap for Disentanglement
$\mathbf{L2^{2}}:c_{\mathcal{X}}=c_{\mathcal{Y}}=\|\cdot-\cdot\|_{2}^{2}$
+ Enc-DST $59.6$$\pm 6.9$ $75.7$$\pm 3.0$ $88.7$$\pm 7.1$ $90.3$$\pm 7.9$
+ Enc-GMG $62.3$$\pm 8.4$ $75.4$$\pm 5.3$ $88.4$$\pm 7.7$ $90.1$$\pm 4.3$
+ Dec-DST $71.5$$\pm 3.6$ $75.8$$\pm 6.6$ $92.1$$\pm 9.7$ $90.9$$\pm 7.6$
+ Dec-GMG $72.0$$\pm 8.5$ $78.9$$\pm 5.0$ $92.5$$\pm 4.4$ $91.7$$\pm 6.0$
$\mathbf{ScL2^{2}}:c_{\mathcal{X}}=\|\cdot-\cdot\|_{2}^{2},c_{\mathcal{Y}}=\alpha\|\cdot-\cdot\|_{2}^{2}$, $\alpha>0$ learnable
+ Jac $61.4$$\pm 12.8$ $76.7$$\pm 4.5$ $90.5$$\pm 3.8$ $91.5$$\pm 5.6$
+ Enc-DST $65.8$$\pm 11.9$ $73.0$$\pm 7.9$ $92.4$$\pm 3.7$ $89.2$$\pm 3.8$
+ Enc-GMG $65.1$$\pm 5.5$ $76.1$$\pm 7.7$ $90.8$$\pm 9.2$ $92.0$$\pm 5.3$
+ Dec-DST $67.4$$\pm 7.1$ $77.9$$\pm 4.5$ $93.2$$\pm 9.7$ $94.5$$\pm 6.9$
+ Dec-GMG 70.0 ±$\pm 5.9$ 81.0 ±$\pm 3.2$ 93.3 ±$\pm 8.6$ 96.1 ±$\pm 3.8$
[…]
+ Enc-DST $69.2$$\pm 9.1$ $77.2$$\pm 7.5$ $87.7$$\pm 7.7$ $90.5$$\pm 5.9$
+ Enc-GMG $70.9$$\pm 9.5$ $79.6$$\pm 6.6$ $92.5$$\pm 5.9$ $93.5$$\pm 6.9$
+ Dec-DST 76.8 ±$\pm 4.1$ 81.3 ±$\pm 4.7$ $87.5$$\pm 3.3$ $91.9$$\pm 9.4$
+ Dec-GMG 82.1 ±$\pm 4.5$ 83.7 ±$\pm 8.8$ 95.7 ±$\pm 5.8$ 96.9 ±$\pm 4.9$
Which costs $c_{\mathcal{X}},c_{\mathcal{Y}}$ should we choose?  The first question that naturally arises when using a geometry-preserving regularizer is: Which geometric features should be preserved? Previous works (Nakagawa et al., 2023, Lee et al., 2022, Huh et al., 2023) focused on preserving scaled distances between points, with the scale being learnable. We follow and extend this approach by also investigating plain distances and angles. This leads to three choices for $c_{\mathcal{X}},c_{\mathcal{Y}}$ ($\mathbf{L2^{2}},\mathbf{ScL2^{2}},\mathbf{Cos}$), as introduced in §3.1, following the hierarchy of geometry-preserving mappings proposed in Lee et al. (2022). We benchmark these on Shapes3D across various settings including DST, GMG, and Jac. Table 1 shows that angle preservation ($\mathbf{Cos}$), previously unconsidered for disentangled representation learning, consistently outperforms (scaled) distance preservation. This result is intuitive, as preserving angles imposes a weaker constraint, allowing for greater latent space expressiveness. In practice, preserving scaled distances seems to overly restrict the expressiveness of the latent space.
Should we regularize the encoder $e_{\phi}$, or the decoder $d_{\theta}$?  The next question we aim to answer is whether the decoder or encoder should be regularized. Therefore, we follow the previous setup on Shapes3D and benchmark all geometry-preserving regularizers on $e_{\phi}$ and $d_{\theta}$ as reported in Table 1. We find that regularizing the decoder is beneficial over regularizing the encoder. We hypothesize this is due to the regularization of $d_{\theta}$ offering a stronger signal as its gradients impact both $\phi$ and $\theta$, as in this case, the reference $r=e_{\phi}\sharp p_{\textrm{data}}$ is the distribution of encoded images. Our findings align with prior works (Lee et al., 2022, Nakagawa et al., 2023), which focus on regularizing $d_{\theta}$ yet do not offer this type of analysis. Additionally, we find that the GMG consistently outperforms the DST over all costs. Overall, the GMG on $d_{\theta}$ with $\mathbf{Cos}$ achieves best DCI-D results over all baselines. Consequently, moving forward we regularize the decoder for angle preservation.
@@
Table 2: Effect of GMG and DST with $\mathbf{Cos}$ as costs on disentanglement, as measured by DCI-D, over three datasets. We highlight the best, and second best result for each dataset and method.
With $\mathbf{Cos}$ costs $\beta$-VAE $\beta$-TCVAE $\beta$-VAE + HFS $\beta$-TCVAE + HFS
DSprites Higgins et al. (2017)